# Conditional Quantile Estimation for Uncertain Watch Time in Short-Video Recommendation

### v1

@@ Abstract.
Within the domain of short video recommendation, predicting users’ watch time is a critical but challenging task. Prevailing deterministic solutions obtain accurate debiased statistical models, yet they neglect the intrinsic uncertainty inherent in user environments. In our observation, we found that this uncertainty could potentially limit these methods’ accuracy in watch-time prediction on our online platform, despite that we have employed numerous features and complex network architectures. Consequently, we believe that a better solution is to model the conditional distribution of this uncertain watch time.
In this paper, we introduce a novel estimation technique—Conditional Quantile Estimation (CQE), which utilizes quantile regression to capture the nuanced distribution of watch time. The learned distribution accounts for the stochastic nature of users, thereby it provides a more accurate and robust estimation. In addition, we also design several strategies to enhance the quantile prediction including conditional expectation, conservative estimation, and dynamic quantile combination. We verify the effectiveness of our method through extensive offline evaluations using public datasets as well as deployment in a real-world video application with over 300 million daily active users.
video recommendation, uncertain watch time, conditional quantile estimation
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems.
@@ 1. Introduction
The rapid growth of online video platforms has significantly changed the way users consume digital content, with short videos emerging as an increasingly popular format (Liu et al., 2021, 2019; Davidson et al., 2010; Zhou et al., 2010). These platforms rely on recommendation systems to curate personalized content that aligns with users’ interests, thereby enhancing user engagement and satisfaction. Different from indicators like click-through-rate and ratings that are widely used in conventional recommendation problems (e.g. e-commerce and news recommendation), a critical metric for measuring the engagement/interest in video recommendation is the watch-time, which is a dense and nuanced signal that reflects the amount of time a user spends on a particular video. Additionally, an accurate watch-time prediction would also help augmenting the ranking models and benefit the video recommendation. Nevertheless, accurately predicting the watch-time presents a formidable challenge due to the complex and dynamic nature of user-video interaction.
Generally, one can formulate the watch-time prediction problem as a regression task. Existing solutions learn a deterministic predictor (Zhan et al., 2022; Tang et al., 2023), which have shown that the watch-time not only depends on the user’s interest but is also related to non-interest factors like the full video’s duration. Despite their effectiveness, the inherent uncertainty in user interactions are not modeled, which potentially limit their accuracy and robustness. For example, a user would watch the entire video of interest in normal cases but may skip the video if he/she is caught up by something else outside the online service. As a solution, some work proposes to model the watch-time as a random variable (Zheng et al., 2022; Zhao et al., 2023; Lin et al., 2023) by assuming a Gaussian, mixture of Gaussian or Categorical distribution, and include a variance estimation that represents how confident the output watch-time is. This approach turns out to be more robust against user’s uncertain behaviors than a deterministic point estimator, and is able to generate more accurate predictions and recommendations. Even though, this strategy inevitably depends on the validity of the pre-defined distribution, which might over-simplify or mismatch the real-world user behavior.
In this paper, we aim to achieve a more nuanced approach to watch-time prediction that overcomes the aforementioned limitations, and introduce our Conditional Quantile Estimation (CQE) framework that explicitly models the watch-time distribution in quantiles. Specifically, it consists of a multi-quantile modeling step that outputs the estimation of evenly separated quantile points for the target watch-time distribution, and a quantile combination step that allows different strategies to integrate the output point estimations. By modeling the watch-time distribution with quantiles, the method becomes agnostic to the underlying watch-time distributions while providing sufficient expressiveness of the user behavior uncertainty. During training, we follow standard quantile regression techniques (Koenker, 2005) known for its robustness and flexibility. After obtaining all quantiles’ estimations, the theoretically most accurate predictions could be achieved with the expectation. Yet, this strategy might not always be suitable when users are not tolerant of bad recommendations or the recommender system requires dynamic controls. In practice, we offer conservative estimation and dynamic quantile combination as alternative strategies to accommodate such scenarios.
In summary, the contributions of this paper are as follows:
  • We introduce the Conditional Quantile Estimation (CQE) framework, a novel estimation method that explicitly models the conditional distribution of watch time.
  • We develop several innovative strategies that extend the CQE model to different recommendation scenarios and user demands.
  • We conduct the comprehensive offline evaluation of the proposed CQE framework, which demonstrates consistent effectiveness in enhancing video consumption.
  • We further verify the superior performance of CQE in a real-world industrial video recommendation setting with billion-scale user interactions.
In the remainder of the paper, we will first illustrate the background of the watch-time prediction task in video recommendation and quantile regression in section 2. Then, we present our solution framework in section 3 and our experiments in section 4.
@@ 2.1. Video Recommendation and Watch-time Prediction
Video recommendation systems have evolved to cater to the growing demand for personalized content delivery. With the advent of online video platforms like YouTube and TikTok, the importance of accurate video recommendation has been underscored by the significant impact on user retention and satisfaction (Liu et al., 2021, 2019; Davidson et al., 2010; Zhou et al., 2010; Tang et al., 2023). In the realm of video recommendation systems, accurately predicting user engagement through watch time is a critical challenge. Watch time serves as a key metric for gauging user interest and engagement with recommended videos. The initial study (Covington et al., 2016) centered on enhancing video recommendations for the YouTube platform, introducing the Weighted Logistic Regression (WLR) technique for predicting watch time. This approach has since been recognized as an advanced method in its field. Nevertheless, the WLR’s applicability is not straightforward to full-screen video recommendation systems, and it could encounter significant bias problems attributable to its weighted calculation system. D2Q (Zhan et al., 2022) mitigates the duration bias through the implementation of backdoor adjustments and modeling the watch-time quantile under different duration groups. $\text{D}^{2}$o (Zhao et al., 2023) tackles the problem of bias in video recommendation watch times by using a model that corrects for both duration bias and noisy watching, providing a more accurate measure of user interest. DVR (Zheng et al., 2022) introduces a novel metric called WTG (Watch Time Gain) and employs adversarial learning to learn unbiased user preferences. Our method can be seamlessly integrated into a wide array of duration debiased methods, thereby significantly improving their predictive accuracy. TPM (Lin et al., 2023) breaks down the task into a series of interconnected classification problems arranged in a tree-like structure. Although TPM considers watch-time variance, it does not capture the full breadth of the watch-time distribution as our method does.
The watch-time prediction task also faces the critical problem of duration bias (Zhao et al., 2023; Zhan et al., 2022; Zheng et al., 2022; Lin et al., 2023). This bias indicates that users prefer to spend more time watching longer-duration videos, skewing the average watch time in favor of longer content. Such a preference for duration over shorter alternatives complicates the task of accurately predicting user engagement. Our methodology integrates seamlessly with most duration debiasing methods, significantly improving their predictive ability.
@@ 2.2. Quantile Regression
Quantile regression is a type of regression analysis widely used in statistics, econometrics, and ecology (Koenker, 2005). Unlike traditional mean/linear regression—which focuses on estimating the average outcome, quantile regression seeks to estimate the conditional median and other quantiles of the random variable. This flexible feature provides a more comprehensive understanding of the variable’s distributional effects that mean regression may overlook(Cade and Noon, 2003; Wei et al., 2006; Wei and He, 2006). In the context of machine learning, quantile regression has been extended beyond linear models. Representative approaches (Rodrigues and Pereira, 2018; Petneházi, 2021; Tang et al., 2022) integrate quantile regression into neural networks, offering means to forecast conditional quantiles in non-linear and high-dimensional settings. QRF (Meinshausen, 2006) further deploys quantile regression within random forests, further exemplifying its adaptability and the enhancement of predictive capabilities across diverse models.
The presented solution in this paper intends to integrate the principles of quantile regression into the domain of video recommendation systems. By adapting this statistical approach to account for the uncertainty and variability in watch time, we propose a novel application that enhances the predictive performance of recommendation systems. This advancement promotes a more nuanced understanding of user engagement, driving toward more personalized and satisfactory user experiences.
@@ 3.1. Problem Formulation
The objective of the recommender system is intrinsically a predictive task, where we aim to estimate the degree of a user’s engagement with the recommended content. In the video recommendation scenario, this engagement is mainly estimated by the user’s watch-time—the time a user spends watching the video. Formally, we represent each user-video pair as $(u,v)$, and assume a pre-defined mapping function that extracts an $n$-dimensional feature vector $\mathbf{x}=\psi(u,v)\in\mathbb{R}^{n}$. These features encompass contextual information, video characteristics, user profile attributes, and the user’s historical interaction data. Considering the user’s watch-time as a random variable $W$, the goal is to estimate the probability distribution of $W$.
3.2. Conditional Quantile Estimation
Figure 1. Proposed conditional quantile estimation framework for watch-time inference.
Previous research simplifies the problem by assuming that the conditional watch-time follows a Gaussian, mixed Gaussian or Categorical distribution  (Zheng et al., 2022; Zhao et al., 2023; Lin et al., 2023). Yet, the user’s behavior might be far more complicated as we have discussed in 1. In contrast, we aim to learn the quantiles of the watch-time without making any assumption about the underlying distribution. Formally, we define a set of $N$ quantiles $\tau_{1},\tau_{2},\dots,\tau_{N}\in(0,1)$ that we are interested in. For example, $\tau=0.5$ corresponds to the median, while $\tau=0.1$ and $\tau=0.9$ correspond to the lower and upper tails of the distribution, respectively. As a special case, when adopting even separation, we have $\tau_{i}=i/(N+1)$. Rather than giving a point estimation of the expected watch-time, we construct a conditional quantile estimation function $\phi(t_{\tau_{1}},\dots,t_{\tau_{N}}|\textbf{x})$ to infer the corresponding watch-time value $t_{\tau_{i}}$ for each of the quantile $\tau_{i}$, given the input $x$. To ensure the values of $\{t_{\tau_{i}}\}_{i=1}^{N}$ are ordered, we will first have the model output $N$ non-negative elements, and then $\{t_{\tau_{i}}\}_{i=1}^{N}$ will be the cumulative sum of these elements. As shown in Figure 1, these output watch-time values exemplify the watch-time distribution, where output quantile values are closer to each other in areas with higher likelihood. This provides richer information about the user behavior, especially for extreme values like very high or very low watch-times.
During training, an ideal solution would sample the watch-time for multiple rounds until sufficient samples are observed for accurate approximation of all $N$ quantiles. However, this is impractical since we cannot require the user to repeatedly watch the same video under identical conditions. To solve this problem, we minimize the pinball loss that penalizes the deviation between the quantile estimates and the actual observed values (Koenker, 2005). For an individual $\tau$, the loss function is defined as
(1) $\mathcal{L}_{\tau}(y,t_{\tau_{i}})=\begin{cases}\tau_{i}(y-t_{\tau_{i}})&\text{if }y\geq t_{\tau_{i}}\\ (1-\tau_{i})(t_{\tau_{i}}-y)&\text{otherwise}\end{cases}$
where $y$ is the actual watch-time. And for each sample $(x,y)$, we minimize the aggregated quantile regression loss:
(2) $\mathcal{L}_{QR}=\sum_{i=1}^{N}\mathcal{L}_{\tau}(y,t_{\tau_{i}}).$
Intuitively, the optimization is slower either when the prediction quantile value $t_{\tau_{i}}$ is close to the ground truth $y$, or its quantile is far from $y$. This would help learning more accurate approximations for quantiles closer to $y$, and avoid misguidance to quantiles on the other side of the distribution.
Without specification, we denote our solution framework as Conditional Quantile Estimation (CQE) for the rest of the paper. Through this method, we can construct a broad spectrum of predictions that allows us to reliably infer the distribution of user watch-time. This more complete understanding of the distribution of watch times potentially improves the precision of the estimation and enhance the performance of recommendation systems as a whole. As we will illustrate in the next section, it also provides a solid foundation for the flexible design of personalized recommendation strategies during inference. Though not the focus of this paper, this estimation can extend to adjust for duration bias or capture user interest inherent in the observed watch-time as in existing works (Zheng et al., 2022; Zhao et al., 2023).
3.3. Inference with Expanded Strategies
Once we have estimated the quantiles of the distribution $t_{\tau_{i},\dots,\tau_{N}}$, we can design a suitable quantile combination strategy for the watch-time inference.

### v4

@@ Abstract.
Accurately predicting watch time is crucial for optimizing recommendations and user experience in short video platforms. However, existing methods that estimate a single average watch time often fail to capture the inherent uncertainty in user engagement patterns. In this paper, we propose Conditional Quantile Estimation (CQE) to model the entire conditional distribution of watch time. Using quantile regression, CQE characterizes the complex watch-time distribution for each user-video pair, providing a flexible and comprehensive approach to understanding user behavior. We further design multiple strategies to combine the quantile estimates, adapting to different recommendation scenarios and user preferences. Extensive offline experiments and online A/B tests demonstrate the superiority of CQE in watch-time prediction and user engagement modeling. Specifically, deploying CQE online on a large-scale platform with hundreds of millions of daily active users has led to substantial gains in key evaluation metrics, including active days, engagement time, and video views. These results highlight the practical impact of our proposed approach in enhancing the user experience and overall performance of the short video recommendation system. The code will be released https://github.com/justopit/CQE.
Video Recommendation, Watch Time Prediction, Conditional Quantile Estimation, Quantile Regression
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
@@ 1. Introduction
The rapid growth of online video platforms has revolutionized the way users consume digital content, and short videos are emerging as one of the most popular formats. Recommender systems play a crucial role in these platforms by providing personalized content recommendations to enhance user engagement and satisfaction (Liu et al., 2021, 2019; Davidson et al., 2010; Zhou et al., 2010; Jing et al., 2024; Quan et al., 2024; Wang et al., 2017, 2021; Yang et al., 2024). Unlike traditional recommendation problems (e.g., e-commerce and news recommendation), a key metric for measuring user interest and engagement in short video recommendation is watch time, which comprehensively reflects users’ preferences and involvement. Therefore, accurately predicting watch time is vital for optimizing recommendation strategies and improving user experience.
However, predicting watch time remains a challenging task due to the inherent uncertainty and heterogeneity in user behavior. In real-world scenarios, it is often infeasible to obtain multiple observations of watch time for the same user-video pair under identical conditions, as users rarely watch the same video multiple times in the exact same context. This limitation prevents us from directly estimating the true conditional watch time distribution from the data.
Existing methods (Zhan et al., 2022; Tang et al., 2023; Zheng et al., 2022; Zhao et al., 2023; Lin et al., 2023; Zhang et al., 2023; Sun et al., 2024) often focus on predicting the conditional expectation of watch time, overlooking the complexity and diversity of the conditional watch time distribution. These methods fail to fully capture the behavioral differences across various user-video pairs, leading to limitations in recommendation performance. The inadequacy of using a single average value to characterize these complex distribution patterns underscores the need to model the entire conditional distribution of watch time.
Figure 1. Illustration of the conditional watch time distribution and some personalized recommendation strategies. (a) Conditional watch time distributions predicted by our CQE model for different user-video pairs, showcasing the heterogeneity and complexity of user engagement patterns. (b) Conservative estimation strategy prioritizing user satisfaction by selecting videos with higher lower quantiles when the expected watch times are similar. (c) Dynamic quantile combination strategy adapting to user churn risk or video novelty, employing lower quantiles for high-churn-risk users or unfamiliar videos, and higher quantiles for low-churn-risk users or familiar videos. (d) Expectation estimation strategy providing a global optimization perspective by considering the entire watch time distribution. We normalize watch time by limiting it to a maximum of 300 seconds and scaling it as min(watch time, 300) / 300.
To address these challenges, we propose the Conditional Quantile Estimation (CQE) model, which predicts the conditional distribution of watch time given a user-video pair and its associated context. As illustrated in Figure 2, CQE leverages quantile regression—a widely adopted method in statistics (Wei et al., 2006a), econometrics (Angrist et al., 2004), reinforcement learning (Dabney et al., 2018), and large language models (Dorka, 2024). By estimating multiple quantiles of the conditional watch time distribution, CQE provides a comprehensive view of potential user engagement patterns. As illustrated in Figure 1(a), the conditional watch time distributions predicted by our CQE model for different user-video pairs exhibit significant diversity in their shapes, peak positions, and dispersion levels. This heterogeneity reflects the intrinsic uncertainty and variability in user preferences and engagement across different contexts.
Capturing the conditional distribution of watch time is essential for understanding user engagement patterns and developing effective recommendation strategies. By considering the detailed characteristics of the watch time distribution, we can gain insight into the diverse watching behaviors of different user groups. This granular understanding enables us to tailor recommendation strategies for various scenarios and user preferences.
Based on the CQE model, we design three primary recommendation strategies. The conservative estimation strategy (Figure 1(b)) prioritizes user satisfaction by selecting videos with higher lower quantiles when the expected watch times are similar, mitigating disengagement risks for users. The dynamic quantile combination strategy (Figure 1(c)) adapts the choice of quantiles based on factors such as the user’s churn risk or video novelty. It assigns more weight to the Low Quantile for high-churn-risk users or novel videos, ensuring a satisfactory experience, and more weight to the High Quantile for low-churn-risk users or familiar videos, potentially offering more engaging recommendations. Finally, the expectation estimation strategy (Figure 1(d)) provides a global optimization perspective, aiming to maximize overall user engagement considering the entire watch time distribution. The diversity of these strategies allows our recommendation system to adapt to different scenarios and user needs, enhancing the quality of personalized recommendations and user experience.
Beyond the heterogeneity shown in Figure 1(a), our empirical analysis on real-world online data quantifies this distributional variability using interquartile range (IQR), defined as $\frac{1}{N/2}\sum_{i=1}^{N/2}t_{\tau_{N-i+1}}-t_{\tau_{i}}$. This analysis reveals a clear relationship between prediction accuracy and distributional spread, with narrower distributions (IQR in $[0.0,0.1]$) achieving significantly higher prediction accuracy (UAUC: $0.721$, MAE: $0.105$) compared to wider distributions (IQR in $[0.5,1.0]$) with poorer performance (UAUC: $0.536$, MAE: $0.305$). This data-driven evidence from production environments strongly supports the need for our conditional quantile approach that adapts to varying degrees of prediction uncertainty.
To this end, we summarize our key contributions in this paper as follows:
  • We propose the CQE model, which employs the quantile regression technique to model the conditional distribution of watch time in short video recommendation, offering a principled approach to capture the uncertainty in user behavior.
  • We design multiple strategies to combine the quantile estimates from CQE, adapting to different recommendation scenarios and user preferences, enhancing the personalization and diversity of the recommendation system.
  • We validate CQE’s superior performance in watch time prediction and user interest modeling through rigorous offline experiments and large-scale online A/B testing, achieving significant improvements in key metrics including active days, engagement duration, and video views.
Figure 2. Illustration of the proposed Conditional Quantile Estimation (CQE) model and its training loss. Left: The CQE model architecture, which takes user, video, and context features as input and outputs multiple quantile estimates of watch time. Right: The pinball loss function used for training the model, showing its asymmetric nature for different quantile levels ($\tau$), allowing the model to learn robust quantile estimates across the entire distribution.
@@ 2.1. Video Recommendation and Watch-time Prediction
Video recommendation systems have evolved to cater to the growing demand for personalized content delivery. With the advent of online video platforms such as YouTube and TikTok, the importance of accurate video recommendation has been underscored by the significant impact on user retention and satisfaction (Liu et al., 2021, 2019; Davidson et al., 2010; Zhou et al., 2010; Tang et al., 2023; Zhang et al., 2023; Zhao et al., 2024). In the realm of video recommendation systems, accurately predicting user engagement through watch time is a critical challenge. Watch time serves as a key metric for gauging user interest and engagement with recommended videos. The initial study (Covington et al., 2016) focused on improving video recommendations for the YouTube platform, introducing the weighted logistic regression (WLR) technique to predict watch time. D2Q (Zhan et al., 2022) mitigates the duration bias through the implementation of backdoor adjustments and modeling the watch-time quantile under different duration groups. $\text{D}^{2}$o (Zhao et al., 2023) uses a model that corrects both duration bias and noisy watching, providing a more accurate measure of user interest. DVR (Zheng et al., 2022) introduces a new metric called WTG (Watch Time Gain) and uses adversarial learning to learn unbiased user preferences.
Unlike existing bucket-based approaches that discretize watch time and focus solely on predicting conditional expectations (e.g., TPM (Lin et al., 2023), CRAED (Sun et al., 2024)), our CQE method employs quantile regression to model the complete continuous distribution. This preserves fine-grained distributional information and captures complex patterns, especially in skewed scenarios where tail behavior is crucial. The continuous nature of CQE enables more flexible inference strategies, including Conservative Estimation and Dynamic Quantile Combination.
Our methodology seamlessly integrates with most duration debiasing techniques, significantly enhancing their predictive performance (Section 4.3.3). Watch-time prediction is fundamentally challenged by duration bias (Zhao et al., 2023; Zhan et al., 2022; Zheng et al., 2022; Lin et al., 2023), where users inherently spend more time on longer videos, skewing average watch time toward extended content. This preference complicates accurate user engagement prediction, making debiasing essential.
@@ 2.2. Quantile Regression
Quantile regression is a type of regression method widely used in statistics (Wei et al., 2006a), econometrics (Angrist et al., 2004), reinforcement learning (Dabney et al., 2018), and large language models (Dorka, 2024). Unlike traditional linear regression, which focuses on estimating the average outcome, quantile regression seeks to estimate the conditional median and other quantiles of the random variable. This flexible feature provides a more comprehensive understanding of the variable’s distributional effects that linear regression may overlook (Cade and Noon, 2003; Wei et al., 2006b; Wei and He, 2006). In the context of machine learning, quantile regression has been extended beyond linear models. Representative approaches (Rodrigues and Pereira, 2018; Petneházi, 2021; Tang et al., 2022) integrate quantile regression into neural networks, offering means to forecast conditional quantiles in nonlinear and high-dimensional settings. QRF (Meinshausen, 2006) further deploys quantile regression within random forests, further exemplifying its adaptability and the enhancement of predictive capabilities across diverse models.
The solution presented in this paper intends to integrate the principles of quantile regression into the domain of video recommendation systems. By adapting this approach to account for the uncertainty and variability in watch time, we propose a novel application that enhances the predictive performance of recommendation systems. This advancement promotes a more nuanced understanding of user engagement, driving toward more personalized and satisfactory user experiences.
@@ 3.1. Problem Formulation
In video recommendation systems, our primary objective is to predict user engagement, typically measured by watch time. Let $(u,v)$ denote a user-video pair under context $c$. We define a feature mapping function $\psi(u,v,c)$ that extracts an $n$-dimensional feature vector $\mathbf{x}\in\mathbb{R}^{n}$. This vector encapsulates user characteristics, video attributes, contextual information, and historical interaction data.
Let $W$ be a random variable representing the watch time. Our goal is to estimate the probability distribution of $W$ conditioned on the input features:
(1) $P(W|\mathbf{x})=P(W|\psi(u,v,c)).$
Unlike traditional approaches (Zhao et al., 2023; Zheng et al., 2022; Covington et al., 2016; Zhan et al., 2022; Tang et al., 2023; Sun et al., 2024; Lin et al., 2023) that focus on estimating the conditional expected watch time $\mathbb{E}[W|\mathbf{x}]$, we aim to characterize the entire conditional distribution. This allows us to capture the inherent uncertainty and variability in user engagement patterns, providing a more comprehensive understanding of potential user behaviors.
3.2. Conditional Quantile Estimation Model
To capture the full distribution of watch time, we propose a Conditional Quantile Estimation (CQE) model. As shown in the left part of Figure 2, this approach allows us to estimate multiple quantiles of the watch time distribution simultaneously, providing a more comprehensive view of potential user engagement.
Let $\{\tau_{1},\tau_{2},...,\tau_{N}\}$ be a set of $N$ predefined quantile levels, where $\tau_{i}=\frac{i}{N+1}$. Our CQE model aims to estimate the corresponding watch time values $\{t_{\tau_{1}},t_{\tau_{2}},...,t_{\tau_{N}}\}$ for each quantile level, given the input features $\mathbf{x}$:
(2) $\{t_{\tau_{1}},t_{\tau_{2}},...,t_{\tau_{N}}\}=\phi(\mathbf{x};\theta),$
where $\phi(\cdot)$ is a neural network parameterized by $\theta$. To ensure the monotonicity of quantile estimates, we implement the following architecture:
(3) $\begin{split}\mathbf{h}&=f(\mathbf{x};\theta_{f}),\\ \mathbf{d}&=\text{ReLU}(g(\mathbf{h};\theta_{g})),\\ t_{\tau_{i}}&=\sum_{j=1}^{i}d_{j},\quad\text{for }i=1,\ldots,N.\end{split}$
Here, $f(\cdot)$ and $g(\cdot)$ are neural network components, $\mathbf{h}$ is an intermediate hidden representation, and $\mathbf{d}$ is a vector of nonnegative elements. The final quantile estimates $t_{\tau_{i}}$ are obtained through cumulative summation, naturally enforcing the ordering constraint $t_{\tau_{1}}\leq t_{\tau_{2}}\leq...\leq t_{\tau_{N}}$.
This formulation allows our model to capture complex, non-linear relationships between input features and conditional watch time quantiles, while maintaining the necessary monotonicity property of quantile functions.
To provide a clear overview of our CQE approach, we present the pseudocode algorithm in Algorithm 1.
Function Train($D,N,\alpha$):
       Input: Training data $D=\{(\mathbf{x}_{i},y_{i})\}$, number of quantiles $N$, learning rate $\alpha$
       Output: Trained CQE model $\phi$
       Initialize model parameters $\theta$;
       Define quantile levels $\tau_{1},...,\tau_{N}$;
       for each epoch do
             for each mini-batch $B\subset D$ do
                   Compute quantile estimates $\{t_{\tau_{1}},...,t_{\tau_{N}}\}=\phi(\mathbf{x};\theta)$;
                   Compute loss $L_{QR}$ using Eq. 5 ;
                   Update $\theta\leftarrow\theta-\alpha\nabla L_{QR}$;
             end for
       end for
      return $\phi$
Function Infer($\phi,\mathbf{x},S$):
       Input: Trained CQE model $\phi$, feature vector $\mathbf{x}$, inference strategy $S$
       Output: Predicted watch time $\hat{y}$
       Compute quantile estimates $\{t_{\tau_{1}},...,t_{\tau_{N}}\}=\phi(\mathbf{x})$;
       Apply inference strategy $S$ to $\{t_{\tau_{1}},...,t_{\tau_{N}}\}$ to get $\hat{y}$;
       return $\hat{y}$
Algorithm 1 Conditional Quantile Estimation (CQE)
The computational complexity of the CQE model is comparable to traditional point estimation methods (Zhao et al., 2023; Zheng et al., 2022), with only a marginal increase due to the estimation of multiple quantiles. In large-scale recommendation systems, the number of unique users and items often reaches hundreds of millions or even billions. These users and items are typically represented by high-dimensional embeddings retrieved using their respective IDs. The number of quantiles required for effective estimation is typically around 100. Therefore, the additional computational cost of CQE is negligible compared to the massive computations required to process the massive user and video features.
3.3. Training Objective
To train our CQE model effectively, we employ the pinball loss function (Niu et al., 2016), which is tailored for quantile regression tasks. For a single quantile level $\tau$, the pinball loss is defined as:
(4) $\mathcal{L}_{\tau}(y,t_{\tau})=\begin{cases}\tau(y-t_{\tau})&\text{if }y\geq t_{\tau};\\ (1-\tau)(t_{\tau}-y)&\text{otherwise};\end{cases}$
where $y$ is the actual watch time and $t_{\tau}$ is the predicted $\tau$-th quantile.
As illustrated in the right part of Figure 2, the pinball loss function has several key properties:
  1. (1)
    Asymmetry: When $\tau\neq 0.5$, the loss is asymmetric around the sample value $y$, with the degree of asymmetry determined by $\tau$.
  2. (2)
    Linearity: The loss increases linearly with the distance between the predicted and actual values, but with different slopes on each side of $y$.
  3. (3)
    Quantile-specific penalties: For $\tau>0.5$, overestimation is penalized more heavily than underestimation, and vice versa for $\tau<0.5$.
These properties make the pinball loss particularly suitable for quantile estimation. For our multi-quantile model, we aggregate the pinball losses across all quantile levels:
(5) $\mathcal{L}_{\text{QR}}=\sum_{i=1}^{N}\mathcal{L}_{\tau_{i}}(y,t_{\tau_{i}}).$
This aggregated loss function encourages the model to learn accurate quantile estimates across the entire distribution, capturing the full spectrum of potential watch times for each user-video pair.
3.4. Inference Strategies
Once we have trained our CQE model to estimate multiple quantiles of the conditional watch time distribution, we can employ various strategies for inference. We propose three main approaches: Conservative Estimation, Dynamic Quantile Combination, and Conditional Expectation. Each strategy offers different advantages and is suitable for specific recommendation scenarios.
3.4.1. Conservative Estimation
In environments where user satisfaction is paramount and the cost of overestimation is high, we adopt a Conservative Estimation (CSE) strategy. This approach focuses on the lower quantiles of the watch time distribution to ensure a satisfactory user experience.
As illustrated in Figure 1(b), when the expected watch times are similar, we prioritize user satisfaction by selecting videos with higher lower quantiles. This strategy helps mitigate the risks of user disengagement.
Formally, with existing expectation estimation in our online system, we select a lower quantile $\tau_{\text{low}}$ (e.g. $\tau_{\text{low}}=0.25$) and use its corresponding watch time prediction:
(6) $\hat{y}_{\text{CSE}}=t_{\tau_{\text{low}}}.$
This strategy helps mitigate the risk of user disappointment due to overly optimistic recommendations, as the actual watch time is likely to exceed this conservative estimate.
3.4.2. Dynamic Quantile Combination
To adapt to varying user preferences and content characteristics, we propose a Dynamic Quantile Combination (DQC) strategy. This approach combines predictions from different quantiles based on contextual factors.
As shown in Figure 1(c), the DQC strategy adapts the choice of quantiles based on the user’s churn risk or video novelty. It assigns more weight to the Low Quantile for high-churn-risk users or novel videos, ensuring a satisfactory experience, and more weight to the High Quantile for low-churn-risk users or familiar videos, potentially offering more engaging recommendations. This dynamic approach allows the system to balance between safe recommendations and potentially more rewarding ones based on the user’s current state and content familiarity.
Let $k\in[0,1]$ be a context-dependent blending parameter, we compute the final prediction as:
(7) $\hat{y}_{\text{DQC}}=k\cdot t_{\tau_{\text{low}}}+(1-k)\cdot t_{\tau_{\text{high}}},$
where $t_{\tau_{\text{low}}}$ and $t_{\tau_{\text{high}}}$ represent conservative and optimistic quantile predictions, respectively.
The blending parameter $k$ can be adjusted based on factors such as user risk profile, video novelty, or platform objectives. For instance, we might use a higher $k$ (favoring conservative estimates) for new users or novel content, and a lower $k$ for established users or familiar content types.